What is the definition of Dipole? Show that the units V/m and N/C are equivalent Why...

  1. What is the definition of Dipole?

  2. Show that the units V/m and N/C are equivalent

  3. Why is there work done when moving a positive charge closer to another positive

    charge?

  4. What are the differences and similarities between the electric field of a point charge

    and a charged sphere?

  5. Consider an electron in an electric field space. Would the electron move at constant

    speed or constant acceleration?
Answer #1

1)A dipole is a pair of equal and opposite charges separated by some distance.

2)from the definition of electric field, its the electric force per unit charge

E = F/q

force has units of N and charge is measured in Coulombs

E = N/C

We know that,

E = V/d = Potential/distance

Potential has units of Volts and distance is measured in m

E = V/m

Hence, V/m and N/C are the units of electric field.

3)Because there exits a potential differnce between them. The work done in moving a charge is given by:

W = q (V2 - V1)

where, q is the charge moving, V2 - V1 is Potential diff between the points.

5)An electron will have constant acceleration.

As in electric field, it will have a force of

F = q E acting on it

F = ma => a = F/m